# Agnostic Smoothed Online Learning

### v2

@@ Abstract
Classical results in statistical learning typically consider two extreme data-generating models: i.i.d. instances from an unknown distribution, or fully adversarial instances, often much more challenging statistically. To bridge the gap between these models, recent work introduced the smoothed framework, in which at each iteration an adversary generates instances from a distribution constrained to have density bounded by $\sigma^{-1}$ compared to some fixed base measure $\mu$. This framework interpolates between the i.i.d. and adversarial cases, depending on the value of $\sigma$. For the classical online prediction problem, most prior results in smoothed online learning rely on the arguably strong assumption that the base measure $\mu$ is known to the learner, contrasting with standard settings in the PAC learning or consistency literature. We consider the general agnostic problem in which the base measure is unknown and values are arbitrary. Along this direction, [15] showed that empirical risk minimization has sublinear regret under the well-specified assumption. We propose an algorithm R-Cover based on recursive coverings which is the first to guarantee sublinear regret for agnostic smoothed online learning without prior knowledge of $\mu$. For classification, we prove that R-Cover has adaptive regret $\tilde{\mathcal{O}}(\sqrt{dT/\sigma})$ for function classes with VC dimension $d$, which is optimal up to logarithmic factors. For regression, we establish that R-Cover has sublinear oblivious regret for function classes with polynomial fat-shattering dimension growth.
@@ 1 Introduction
On one extreme, one can consider that the sequence $(x_{t})_{t\geq 1}$ is fully adversarial and may depend on the actions of the learner. In this case, classical results [38, 5] show that the best one can hope for is to achieve low excess loss compared to function classes with finite Littlestone dimension. This is quite restrictive, for instance, this precludes positive results even for the simple function class of threshold functions $x\mapsto\mathbbm{1}_{x\geq x_{0}}$ on $\mathcal{X}=[0,1]$.
On the other hand, one can suppose that the instance sequence $(x_{t})_{t\geq 1}$ is i.i.d. typically under some unknown distribution $\mu$. In the PAC learning setting [49, 50, 48], one can instead ensure low excess error compared to function classes with finite VC dimension (see Definition 2) which is significantly weaker than having finite Littlestone dimension. For instance, this covers the class of linear separators for say $\mathcal{X}=\mathbb{R}^{d}$ for $d\geq 1$. In regression, this can be replaced with the notion of fat-shattering dimension (see Definition 3) [13, 37], which is a scale-dependent version of the VC dimension. In fact, when the data generating process is i.i.d. one can achieve consistency—vanishing average excess loss—without further function class assumptions [^1] Note that this differs from the PAC learning setting in the sense that guarantees are asymptotic.. For instance, in classification and when the instance space $\mathcal{X}$ is Euclidean, the simple $k$-nearest neighbor algorithm is already consistent [21, 22, 27] under reasonable choices of $k(t)$. Similar consistency results can also be achieved for general spaces [34, 29].
Ideally, one would aim to obtain similar guarantees as for the more amenable i.i.d. case under weaker statistical assumptions. Notably, there has been significant work to establish consistency results under non-i.i.d. instance processes $(x_{t})_{t\geq 1}$, including relaxations of the i.i.d. assumption such as stationary ergodic processes [42, 28, 27] or processes satisfying some form of law of large numbers [26, 47]. More recently, [31] initiated a line of work on universal learning to characterize minimal assumptions on instance processes $(x_{t})_{t\geq 1}$ for consistency [3, 10, 9, 7, 6]. These consistency results are however mostly asymptotic in nature.
@@
Crucially, the above-mentioned works on the standard smooth online learning problem assume that the base measure $\mu$ is known to the learner. Arguably, this is a somewhat strong assumption both in practice and in theory. Knowing the base measure significantly diverges from classical results in the PAC learning setting for which knowing the distribution of the data is unnecessary, or from results from the literature on consistency which require no prior knowledge on the data-generating process. Hence, we aim to answer the following question:
@@
Our contributions.
We answer positively to the previous question by providing a proper algorithm R-Cover (Recursive Covering) based on recursive coverings that achieves the optimal regret guarantee in classification for function classes $\mathcal{F}$ with finite VC dimension up to logarithmic factors (Theorem 4), and sublinear regret in regression for function classes with standard fat-shattering dimension growth (Theorem 6). To the best of our knowledge, this is the first algorithm with sublinear regret guarantees for the general agnostic online learning problem without prior knowledge of the base measure. We note that R-Cover also does not require the knowledge of the smoothness parameter $\sigma$.
@@
where $\tilde{\mathcal{O}}$ hides poly-logarithmic factors in $T$ only. This matches a lower bound for VC classes up to logarithmic factors concurrently obtained by the authors from [15] (confirmed via personal communication). In particular, R-Cover has optimal dependency in $T$, $d$, but also the smoothness parameter $\sigma$. Precisely, there is a function class of VC dimension $d$ for which any learning algorithm must incur an expected regret $\sqrt{dT/\sigma}$ for some smooth adversary (Theorem 5). This lower bound holds even in the realizable setting (well-specified and noiseless) in which there exists some function $f^{\star}\in\mathcal{F}$ fixed a priori for which $y_{t}=f^{\star}(x_{t})$ for all $t\geq 1$, and the loss is fixed over time.
The proof of the regret guarantees of R-Cover crucially relies on a novel property that we prove for smooth adversaries (see Propositions 9 and 13). At the high level, this tightly bounds the possible amount of exploration of unknown regions of the instance space for smooth adversaries. This may be of broader interest for smoothed analysis without prior knowledge of the base measure, or for understanding which relaxations of the smoothness assumption could be made while preserving regret guarantees.
@@ 2 Preliminaries
2.1 Formal setup
We start by formally defining the online learning problem. Let $\mathcal{X}$ be an instance space equipped with some sigma-algebra. The function class $\mathcal{F}$ is a set of measurable functions $f:\mathcal{X}\to[0,1]$. We fix a horizon $T\geq 1$ and consider the following sequential prediction task. At each iteration $t\in[T]$,
@@ 2.1 Formal setup
In particular, this captures the standard prediction setting in which there is a fixed $1$-Lipschitz loss $\ell:[0,1]\times[0,1]\to[0,1]$ and the loss of the learner is given as $\ell(\hat{y}_{t},y_{t})$ for some value $y_{t}$ that is revealed after the prediction $\hat{y}_{t}$. Indeed, the adversary may choose the loss $\ell_{t}(\cdot)=\ell(\cdot,y_{t})$ in step $1$. Next, we say that the learner is proper if at each iteration $t\in[T]$, before observing the query $x_{t}$, the learner first commits to a function $\hat{f}_{t}\in\mathcal{F}$ then, upon observing $x_{t}$, predicts the value $\hat{y}_{t}=\hat{f}_{t}(x_{t})$. Our proposed algorithms will enjoy this property.
The smoothness assumption constrains the choices for the distributions $\mu_{t}$ chosen by the adversary as defined below.
@@
in which the benchmark function may depend on the specific realizations of the learning process, and the expected oblivious regret
@@
in which the benchmark function is fixed prior to the learning process. Adaptive benchmark are known to require significantly stronger analysis than oblivious benchmarks for smoothed online learning (e.g. see [36]).
2.2 Complexity notions for the function class and prior results
In classification, when the functions take value in $\{0,1\}$, when the instance process is i.i.d. ($\sigma=0$) it is known that in our setup, learnability is characterized by the VC dimension [49, 50, 48].
@@ 2.2 Complexity notions for the function class and prior results
Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class for classification. We say that $\mathcal{F}$ shatters a set of points $\{x_{1},\ldots,x_{m}\}\subset\mathcal{X}$ if for any choice of values $\epsilon\in\{0,1\}^{m}$ there exists $f_{\epsilon}\in\mathcal{F}$ such that $f_{\epsilon}(x_{i})=\epsilon_{i}$ for all $i\in[m]$. The VC dimension of $\mathcal{F}$ is the size of the largest shattered set.
We next state the classical Sauer-Shelah’s lemma [45, 46] which bounds the size of the projection of a function class with finite VC dimension onto a set $\{x_{1},\ldots,x_{n}\}\subset\mathcal{X}$.
Lemma 1 (Sauer-Shelah’s lemma).
@@
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class for regression. Fix a scale $\alpha>0$. We say that $\mathcal{F}$ $\alpha$-shatters a set of points $\{x_{1},\ldots,x_{m}\}\subset\mathcal{X}$ if there exist values $s_{1},\ldots,s_{m}\in[0,1]$ such that for any choice of signs $\epsilon\in\{\pm 1\}^{m}$ there exists $f_{\epsilon}\in\mathcal{F}$ such that $\epsilon(f_{\epsilon}(x_{i})-s_{i})\geq\alpha$ for all $i\in[m]$. The fat-shattering dimension of $\mathcal{F}$ at scale $\alpha>0$, denoted $\mathsf{fat}_{\mathcal{F}}(\alpha)$ is the size of the largest $\alpha$-shattered set.
Generalizing Sauer-Shelah’s lemma, it is known that the fat-shattering dimension can be used to bound the size of empirical covers regression function classes. Before stating the bound, we formally define the notion of covering set and covering numbers. We voluntarily restrict these notions to the empirical infinite norm, which is sufficient for this work.
@@
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class for regression. Fix a set $S=\{x_{1},\ldots,x_{n}\}\subset\mathcal{X}$ and $\epsilon\geq 0$. We say that $\mathcal{C}\subset\mathcal{F}$ is an $\epsilon$-cover of $\mathcal{F}$ on $S$ if for all $f\in\mathcal{F}$ there exists $g\in\mathcal{C}$ such that
$\max_{i\in[n]}|f(x_{i})-g(x_{i})|\leq\epsilon.$
The $\epsilon$-covering number of $\mathcal{F}$ on $S$, denoted $\mathcal{N}(\mathcal{F};\epsilon,S)$ is the size of the smallest $\epsilon$-cover of $\mathcal{F}$ on $S$.
The following result bounds these covering numbers similarly to Sauer-Shelah’s Lemma 1.
Theorem 2 (Theorem 4.4 from [44]).
To state some of our results, we also need to define the Wills functional [53, 30] of $\mathcal{F}$, which was first introduced in the context of lattice point enumeration. The definition below uses the formulation from [15].
@@
Note that the above definition depends on the choice of $Z_{1},\ldots,Z_{m}$. For simplicity we may omit this dependency—most of the time we will take its expectation for $Z_{1},\ldots,Z_{m}\overset{iid}{\sim}\mu$. The properties of the Wills functional have been extensively studied [53, 30, 39, 41]. We refer to [41] for detailed connections with metric complexities and universal coding. We give in Appendix A a brief overview of links between Wills functional and other more standard measures complexities that are most relevant to this work, including the VC and fat-shattering dimensions.
In particular, we have $\ln W_{m}(\mathcal{F})\lesssim d\ln m$ for classes $\mathcal{F}$ with finite VC dimension. [41] also showed that we can bound $\ln W_{m}(\mathcal{F})\leq\mathcal{G}_{m}(\mathcal{F})$ where $\mathcal{G}_{m}(\mathcal{F})$ is the Gaussian complexity of $\mathcal{F}$ (see Appendix A for a definition). Last, [15] showed that having $\ln W_{m}(\mathcal{F})=o(m)$ is necessary and sufficient to ensure learnability with polynomially many samples when the data is i.i.d.
Now that we have defined the Wills functional, we can formally state the main result from [15] which shows that empirical risk minimization (ERM) achieves sublinear regret without knowledge of the base measure for the specific well-specified setting.
@@
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function a function class. Consider the squared loss regression setting in which $\ell_{t}(\cdot)=(\cdot-y_{t})^{2}$ for a value $y_{t}\in\mathbb{R}$. Suppose that there exists some function $f^{\star}\in\mathcal{F}$ such that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ and that the values are given via $y_{t}=f^{\star}(x_{t})+\eta_{t}$ where $\eta_{t}\mid\mathcal{H}_{t-1}$ is a mean-zero subgaussian random variable with variance proxy $\nu^{2}$. Then, ERM makes predictions $\hat{y}_{t}$ such that
@@
2.3 Further definitions and notations
@@
Let $(Z_{t})_{t\geq 1}$ be a sequence of random variables adapted to a filtration $(\mathcal{F}_{t})_{t\geq 1}$. A tangent sequence $(Z_{t}^{\prime})_{t\geq 1}$ is a sequence of random variables such that $Z_{t}$ and $Z_{t}^{\prime}$ are i.i.d. conditionally on $\mathcal{F}_{t-1}$.
@@ 3 Main results
We start by giving our main regret guarantees for our algorithm R-Cover in Section 3.1. We then construct in Section 3.2 the algorithm R-Cover instantiated for classification in which case $\mathcal{F}$ is a function class with finite VC dimension $d$. The classification case already provides most of the necessary intuitions and for ease of presentation, we defer the construction of the algorithm in the general regression case to Section 5.1. Last, R-Cover requires a specific variant for a learning with expert advice algorithm which is defined in Section 3.3.
3.1 Main regret guarantees
While our analysis provides regret bounds for general regression function classes, these are more easily stated for classification. In particular, we obtain the following result.
@@ 3.1 Main regret guarantees
As a by-product of the analysis, we also provide a high-probability version of the above expected adaptive regret bound (see Eq. 30). Note that compared to the regret bound Theorem 3 which becomes $\sigma^{-1}\ln^{7/2}(T)\sqrt{dT}$ for VC classes, our regret bound holds for adversarial values $(y_{t})_{t\in[T]}$ and has an improved dependency in $\sigma$: it grows as $1/\sqrt{\sigma}$ instead of $1/\sigma$. In particular, this yields non-trivial regret bounds for any $\sigma\in[d/T,1]$. Our regret bound for R-Cover is complemented by a matching lower bound up to logarithmic factors, which holds even in the realizable noiseless setting. Confirmed by personal communication, the authors from [15] generalized their lower bound for the regret empirical risk minimization (ERM) (Theorem 3) to general algorithms for VC classes, leading to the same result as below. We include the proof in Appendix C for completeness. The proof strategy is also of independent interest and can be used to show that some of the properties we develop on smooth adversaries (Propositions 9 and 13) are essentially tight. We refer to Section 4.1 for further discussion.
@@
As a remark, R-Cover uses a somewhat complex recursive construction to achieve the optimal regret guarantee from Theorem 4. Achieving (worse) sublinear regret without prior knowledge of the base measure with a simpler algorithm is nevertheless possible. In Section 4.1 we describe a very simple and intuitive Cover algorithm which essentially corresponds to the single-depth version of R-Cover and for instance enjoys a $\approx T^{2/3}$ regret guarantee in classification. We refer to Section 4.1 for details on this result which may be of independent interest. Obtaining regret guarantees in regression for Cover is also possible with the same tools developed for R-Cover and we omit details for simplicity.
We next turn to the general regression setting. At the high level, our algorithm for classification is generalized to regression by constructing $\epsilon$-coverings of the function class for some scale $\epsilon$ that is used as a parameter (for VC classes we simply use $\epsilon=0$). In practice, the optimal choice of the scale $\epsilon$ lies in $[1/T,1]$ and only depends on the growth of the fat-shattering dimensions of $\mathcal{F}$. We note, however, that tuning this parameter $\epsilon$ can be fully side-stepped by performing any learning with expert advice algorithm using as experts the algorithms R-Cover for different choice of parameters $\epsilon\in\{2^{-l},l\leq\log_{2}T\}$. The resulting algorithm would enjoy the same regret guarantees as for the optimally-tuned algorithm.
The full version of our regret bound is stated in Theorem 14. For readability, we instantiate the bound for standard growth scenarios for the fat-shattering dimension of $\mathcal{F}$.
@@
Fix $T\geq 1$. Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class. and suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. There exists a universal constant $C>0$ such that we have the following bounds on the oblivious regret of R-Cover, where we denote by $\hat{y}_{t}$ the predictions of the algorithm.
@@
As for Theorem 4, our analysis also provides high-probability versions of the bounds in Theorem 6 (see Theorem 14). Note that the guarantee for classification from Theorem 4 bounds the expected adaptive regret, while in the regression case, Theorem 6 bounds the expected oblivious regret. We leave open the question of whether one can achieve guarantees for the adaptive regret in this case.

### v3

@@ Abstract
Classical results in statistical learning typically consider two extreme data-generating models: i.i.d. instances from an unknown distribution, or fully adversarial instances, often much more challenging statistically. To bridge the gap between these models, recent work introduced the smoothed framework, in which at each iteration an adversary generates an instance from a distribution constrained to have density bounded by $\sigma^{-1}$ compared to some fixed base measure $\mu$. This framework interpolates between the i.i.d. and adversarial cases, depending on the value of $\sigma$. For the classical online prediction problem, most prior results in smoothed online learning rely on the arguably strong assumption that the base measure $\mu$ is known to the learner, contrasting with standard settings in the PAC learning or consistency literature. We consider the general agnostic problem in which the base measure is unknown and values are arbitrary. In this direction, [15] showed that empirical risk minimization has sublinear regret under the well-specified assumption. We propose an algorithm R-Cover based on recursive coverings which is the first to guarantee sublinear regret for agnostic smoothed online learning without prior knowledge of $\mu$ and without the well-specified assumption. For classification, we prove that R-Cover has adaptive regret $\tilde{\mathcal{O}}(\sqrt{dT/\sigma})$ for function classes with VC dimension $d$, which is optimal up to logarithmic factors. For regression, we establish that R-Cover has sublinear oblivious regret for function classes with polynomial fat-shattering dimension growth.
@@ 1 Introduction
On one extreme, one can consider that the sequence $(x_{t})_{t\geq 1}$ is fully adversarial and may depend on the actions of the learner. In this case, classical results [38, 5] show that the best one can hope for is to achieve low excess loss compared to function classes with finite Littlestone dimension. This is quite restrictive, for instance, this precludes positive results even for the simple function class of threshold functions $\{x\in[0,1]\mapsto\mathbbm{1}_{x\geq x_{0}},x_{0}\in[0,1]\}$ on $\mathcal{X}=[0,1]$.
On the other extreme, one can suppose that the instance sequence $(x_{t})_{t\geq 1}$ is i.i.d. typically under some unknown distribution $\mu$. In the PAC learning setting [49, 50, 48], one can ensure low excess error compared to function classes with finite VC dimension (see Definition 2) which is significantly weaker than having finite Littlestone dimension. For instance, this covers the class of linear separators for say $\mathcal{X}=\mathbb{R}^{d}$ for $d\geq 1$. In regression, this can be replaced with the notion of fat-shattering dimension (see Definition 3) [13, 37], which is a scale-dependent version of the VC dimension. In fact, when the data generating process is i.i.d. one can achieve consistency—vanishing average excess loss—without further function class assumptions [^1] Note that this differs from the PAC learning setting in the sense that guarantees are asymptotic.. For instance, in classification and when the instance space $\mathcal{X}$ is Euclidean, the simple $k$-nearest neighbor algorithm is already consistent [21, 22, 27] under reasonable choices of $k(t)$. Similar consistency results can also be achieved for general spaces [34, 29].
Ideally, one would aim to obtain similar guarantees as for the more amenable i.i.d. case under weaker statistical assumptions. Notably, there has been significant work to establish consistency results under non-i.i.d. instance processes $(x_{t})_{t\geq 1}$, including relaxations of the i.i.d. assumption such as stationary ergodic processes [42, 28, 27] or processes satisfying some form of law of large numbers [26, 47]. More recently, [31] initiated a line of work on universal learning to characterize minimal assumptions on instance processes $(x_{t})_{t\geq 1}$ for consistency [3, 10, 9, 7, 6]. These results are however mostly asymptotic in nature.
@@
Crucially, the above-mentioned works on the standard smooth online learning problem assume that the base measure $\mu$ is known to the learner. Arguably, this is a somewhat strong assumption both in practice and in theory. Knowing the base measure significantly diverges from classical results in the PAC learning setting for which knowing the distribution of the data is unnecessary, or from results from the literature on consistency which require no prior knowledge on the data-generating process. Hence, we aim to answer the following:
@@
Contributions.
We answer positively to the previous question by providing a proper algorithm R-Cover (Recursive Covering) that achieves the optimal regret guarantee in classification for function classes $\mathcal{F}$ with finite VC dimension up to logarithmic factors (Theorem 2), and sublinear regret in regression for function classes with standard fat-shattering dimension growth (Theorem 4). To the best of our knowledge, this is the first algorithm with sublinear regret guarantees for the general agnostic online learning problem without prior knowledge of the base measure. R-Cover also does not require the knowledge of the smoothness parameter $\sigma$.
@@
where $\tilde{\mathcal{O}}$ hides poly-logarithmic factors in $T$ only. This matches a lower bound for VC classes up to logarithmic factors concurrently obtained by the authors from [15]. In particular, R-Cover has optimal dependency in $T$, $d$, but also in the smoothness parameter $\sigma$. More precisely, there is a function class of VC dimension $d$ for which any learning algorithm must incur an expected regret $\sqrt{dT/\sigma}$ for some smooth adversary (Theorem 3). This lower bound holds even in the realizable setting (well-specified and noiseless) in which there exists some function $f^{\star}\in\mathcal{F}$ fixed a priori for which $y_{t}=f^{\star}(x_{t})$ for all $t\geq 1$, and the loss is fixed over time.
The proof of the regret guarantees of R-Cover crucially relies on a novel property that we prove for smooth adversaries (see Propositions 7 and 11). At a high level, this tightly bounds the possible amount of exploration of unknown regions of the instance space for smooth adversaries, which may be of broader interest.
@@ 2 Preliminaries
Formal setup.
Let $\mathcal{X}$ be an instance space equipped with some sigma-algebra. The function class $\mathcal{F}$ is a set of measurable functions $f:\mathcal{X}\to[0,1]$. We fix a horizon $T\geq 1$ and consider the following sequential prediction task. At each iteration $t\in[T]$,
@@
This captures the standard online prediction problem in which there is a fixed $1$-Lipschitz loss $\ell:[0,1]\times[0,1]\to[0,1]$ and the loss of the learner is equal to $\ell(\hat{y}_{t},y_{t})$ for some value $y_{t}$ that is revealed after the prediction $\hat{y}_{t}$. Next, we say that the learner is proper if at each iteration $t\in[T]$, before observing the query $x_{t}$, the learner first commits to a function $\hat{f}_{t}\in\mathcal{F}$ then, upon observing $x_{t}$, predicts the value $\hat{y}_{t}=\hat{f}_{t}(x_{t})$. Our proposed algorithms will enjoy this property.
The smoothness assumption constrains the distributions $\mu_{t}$ chosen by the adversary.
@@
in which the benchmark function may depend on the specific realizations of the learning process, and the expected oblivious regret in which the benchmark function is fixed a priori:
@@
Adaptive benchmarks are known to require significantly stronger analysis than oblivious benchmarks for smoothed online learning (e.g. see [36]).
Complexity notions for the function class and prior results.
In classification, i.e., when the functions take value in $\{0,1\}$, and when the instance process is i.i.d. ($\sigma=0$) it is known that in our setup, learnability is characterized by the VC dimension [49, 50, 48].
@@
Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class. We say that $\mathcal{F}$ shatters a set of points $\{x_{1},\ldots,x_{m}\}\subset\mathcal{X}$ if for any choice of values $\epsilon\in\{0,1\}^{m}$ there exists $f_{\epsilon}\in\mathcal{F}$ such that $f_{\epsilon}(x_{i})=\epsilon_{i}$ for all $i\in[m]$. The VC dimension of $\mathcal{F}$ is the size of the largest shattered set.
@@
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class. Fix $r>0$. We say that $\mathcal{F}$ $r$-shatters a set $\{x_{1},\ldots,x_{m}\}\subset\mathcal{X}$ if there exist $s_{1},\ldots,s_{m}\in[0,1]$ such that for any signs $\epsilon\in\{\pm 1\}^{m}$ there exists $f_{\epsilon}\in\mathcal{F}$ such that $\epsilon(f_{\epsilon}(x_{i})-s_{i})\geq r$ for all $i\in[m]$. The fat-shattering dimension of $\mathcal{F}$ at scale $r>0$, denoted $\mathsf{fat}_{\mathcal{F}}(r)$ is the size of the largest $r$-shattered set.
Next, we define the notion of covering set and covering numbers. We voluntarily restrict these notions to the empirical infinite norm, which is sufficient for this work.
@@
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class for regression. Fix a set $S=\{x_{1},\ldots,x_{n}\}\subset\mathcal{X}$ and $\epsilon\geq 0$. We say that $\mathcal{C}\subset\mathcal{F}$ is an $\epsilon$-cover of $\mathcal{F}$ on $S$ if for all $f\in\mathcal{F}$ there exists $g\in\mathcal{C}$ such that for all $i\in[n]$, $|f(x_{i})-g(x_{i})|\leq\epsilon.$ The $\epsilon$-covering number of $\mathcal{F}$ on $S$, denoted $\mathcal{N}(\mathcal{F};\epsilon,S)$ is the size of the smallest $\epsilon$-cover of $\mathcal{F}$ on $S$.
To state related results, we also need to define the Wills functional [53, 30] of $\mathcal{F}$, which is a less standard complexity measure. The definition below uses the formulation from [15].
@@
The above definition depends on the choice of $Z_{1},\ldots,Z_{m}$. For simplicity we may omit this dependency—most of the time we take its expectation for $Z_{1},\ldots,Z_{m}\overset{iid}{\sim}\mu$. Properties of the Wills functional have been extensively studied [53, 30, 39, 41]. We refer to [41] for detailed connections with metric complexities and universal coding. Appendix A gives a brief overview of links between Wills functional and more standard measures complexities that are most relevant to this work. In particular, for any choice of $Z_{1},\ldots,Z_{m}$, $\ln W_{m}(\mathcal{F})\lesssim d\ln m$ for classes $\mathcal{F}$ with finite VC dimension. [41] also showed that $\ln W_{m}(\mathcal{F})\leq\mathcal{G}_{m}(\mathcal{F})$ where $\mathcal{G}_{m}(\mathcal{F})$ is the Gaussian complexity of $\mathcal{F}$ (see Appendix A for a definition). Last, [15] showed that having $\ln W_{m}(\mathcal{F})=o(m)$ is necessary and sufficient to ensure learnability with polynomially many samples when the data is i.i.d.
Now that we have defined the Wills functional, we can formally state the main result from [15] which shows that empirical risk minimization (ERM) achieves sublinear regret without knowledge of the base measure for the well-specified setting.
@@
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function a function class. Consider the squared loss regression setting in which $\ell_{t}(\cdot)=(\cdot-y_{t})^{2}$ for a value $y_{t}\in\mathbb{R}$. Suppose that there exists some function $f^{\star}\in\mathcal{F}$ such that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ and that the values are given via $y_{t}=f^{\star}(x_{t})+\eta_{t}$ where $\eta_{t}$ is a mean-zero subgaussian random variable with variance proxy $\nu^{2}$, conditionally on the history up to time $t$. Then, ERM makes predictions $\hat{y}_{t}$ such that
@@
Further definitions and notations.
@@
Let $(Z_{t})_{t\geq 1}$ be a sequence of random variables adapted to a filtration $(\mathcal{F}_{t})_{t\geq 1}$. A tangent sequence $(Z_{t}^{\prime})_{t\geq 1}$ is a sequence of random variables such that $Z_{t}$ and $Z_{t}^{\prime}$ are i.i.d. conditionally on $\mathcal{F}_{t-1}$ (and independently of $Z_{t^{\prime}}$ for $t^{\prime}>t$).
@@ 3 Main results
While our analysis provides regret bounds for general regression function classes, these are more easily stated for classification.
@@
As a by-product of the analysis, we also provide a high-probability version of the above regret bound (see Eq. 31). Compared to the regret bound Theorem 1 which becomes $\sigma^{-1}\ln^{7/2}(T)\sqrt{dT}$ for VC classes, our regret bound holds for adversarial values $(y_{t})_{t\in[T]}$ and has an improved dependency in $\sigma$: it grows as $1/\sqrt{\sigma}$ instead of $1/\sigma$. Our regret bound for R-Cover is complemented by a matching lower bound up to logarithmic factors, which holds even in the realizable noiseless setting. The authors from [15] also generalized their lower bound for the regret empirical risk minimization (ERM) (Theorem 3) to general algorithms for VC classes, leading to the same result as below. We include the proof in Appendix C for completeness.
@@
R-Cover uses a somewhat complex recursive construction to achieve the optimal regret guarantee from Theorem 2. For intuition, we also describe in Section 4.1 a very simple and intuitive algorithm Cover that achieves (worse) sublinear regret without prior knowledge of the base measure. This algorithm which essentially corresponds to the single-depth version of R-Cover and for instance enjoys a $\approx T^{2/3}$ regret guarantee in classification, which may be of independent interest. Obtaining regret guarantees in regression for Cover is also possible with the same tools developed for R-Cover and we omit details for simplicity.
We next turn to the general regression setting. At the high level, our algorithm for classification is generalized to regression by constructing $\epsilon$-coverings of the function class for some scale $\epsilon$ that is used as a parameter (for VC classes we simply use $\epsilon=0$). In practice, the optimal choice of the scale $\epsilon$ lies in $[1/T,1]$ and only depends on the growth of the fat-shattering dimensions of $\mathcal{F}$. However, tuning this parameter $\epsilon$ can be fully side-stepped by performing any learning with expert advice algorithm using as experts the algorithms R-Cover for different choice of parameters $\epsilon\in\{2^{-l},l\leq\log_{2}T\}$. The resulting algorithm would enjoy the same regret guarantees as for the optimally-tuned algorithm. The full version of our regret bound is stated in Theorem 12. For readability, we instantiate the bound for standard growth scenarios of the fat-shattering dimension.
@@
Fix $T\geq 1$. Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class and suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. There exists a universal constant $C>0$ such that we have the following bounds on the oblivious regret of R-Cover, where we denote by $\hat{y}_{t}$ the predictions of the algorithm.
@@
Our analysis also provides high-probability versions of these bounds (see Theorem 12). Note that the guarantee for classification from Theorem 2 bounds the expected adaptive regret, while in the regression case, Theorem 4 bounds the expected oblivious regret. We leave open the question of whether one can achieve guarantees for the adaptive regret in regression.
In the rest of this section, we first construct in Section 3.1 the algorithm R-Cover instantiated for classification. This already provides most of the necessary intuitions and for ease of presentation, we defer the construction of the algorithm in the general regression case to Section 5.1. R-Cover requires a specific variant for a learning with expert advice algorithm defined in Section 3.2.